# Simulated optical molasses cooling of trapped antihydrogen

### v2

@@ I Introduction
The most promising cooling technique, and the only one successfully implemented, is a simple optical molasses based on the 1S-2P transition. For this case, a single 121.6 nm beam, somewhat red detuned from the transition in the magnetic field, cools the axial motion of the $\bar{\rm H}$ atoms. For this geometry, the random re-emission of the photon tends to heat the radial motion. Because the magnetic trapping potentials are not perfectly symmetric, the $\bar{\rm H}$ motion mixes the axial and radial degrees of freedomSurkov et al. (1994); Zhong et al. (2018) which can lead to cooling of all 3 directions.Donnan et al. (2013); Baker et al (ALPHA collaboration) The intensity of the laser has to be sufficiently low that there is time for the mixing to take place between successive photon scattering otherwise the atoms will heat on average. This heating is strongest for smallest detuning where the photon scattering rate is largest. The spatially varying magnetic field also complicates the photoabsorption by giving substantial changes in the detuning versus position in the trap.
@@ II Energy levels in magnetic field
In the ALPHA experiment, $\bar{\rm H}$s are laser cooled in a magnetic trap with a minimum $B\sim 1$ TBaker et al (ALPHA collaboration). This means the energy levels are strongly changed from their low (or zero) field character. Reference Rasmussen et al. (2017) discusses the energy levels. The energies in a magnetic field are labeled with a Roman letter at the end which, by convention, increases from low to high energy for 1S states while increasing from high to low energy for the 2P states (as ordered for small $B$). There are two 1S states that can be trapped called the 1Sc and 1Sd states. The 1Sd state has the positron and antiproton spin aligned; in the 1Sc state, they are antiparallel. To prevent losses from the photon emission step, the laser will excite the 2Pa state which has the positron spin and orbital angular momentum aligned to give total positron angular momentum $J=3/2,M=3/2$. Because of the different magnetic moments, the transition frequency for the 1Sc-2Pa transition is approximately 675 MHz higher than that for the 1Sd-2Pa transition at 1 T. The difference in frequencies does not change much with increasing $B$; for example, at $B=$0.5 T, the difference in frequencies is 660 MHz while the difference in frequencies is 680 MHz at 1.5 T.
@@
Figure 1: The splittings of the 1S and 2P states relative to their zero field average. The hyperfine terms of the 2P states are too small to resolve.
@@ III.2 Ideal Optical Molasses
In this section, we describe a method for Monte Carlo simulations of laser cooling of $\bar{\rm H}$ using an ideal optical molassesFoot (2005). We use a method similar to Sec. IV but is idealized by having the velocity mix much faster than the times between photon scattering and by not including the shift in detuning as the $\bar{\rm H}$ moves through the trap. Ideal optical molasses cooling assumes the photon has a wave vector, $\vec{k}$, in a random direction relative to the atom’s velocity, $\vec{v}$. The Doppler effect leads to an effective detuning of
@@
Our Monte Carlo simulation of the ideal optical molasses used the following algorithm: (1) for each of $N$ atoms, randomize the direction of the atom’s velocity, $\vec{v}$, and for that atom compute the effective detuning, Eq. (4); (2) compute the probability that the photon was absorbed using $P=P_{0}/[1+(2\Delta/\Gamma)^{2}]$ with $P_{0}$ a small number of order 0.01; (3) compute a random number; if it is smaller than $P$, then a photon was absorbed; (4) if a photon is absorbed, compute a random emission direction and use Eq. (5) to update that atom’s velocity. If we wanted to simulate the effect of thermalization from elastic collisions between the atoms, we added a step: (5) randomly pick pairs of atoms and rotate their relative velocity to emulate a collision; repeated often enough step (5) leads to a Maxwell-Boltzmann velocity distribution at temperature $k_{B}T=(2/3)\langle KE\rangle$.
@@
The minimum temperature that can be reached for $1M$ happens for a detuning of $\Delta_{0}\simeq-1.1\Gamma$ and gives $T\simeq 3.05$ mK. The temperature for $1M$ at $\Delta_{0}=-\Gamma/2$ (the minimum of Eq. (6)) is $\simeq 19$ mK which is almost 8×8$\times$8 × hotter than expected from the usual optical molasses relations. From Fig. 2(a), the higher masses become progressively closer to the ideal optical molasses case, red solid line. The reason for the discrepancy is the large size of the velocity kick, 3.26 m/s, compared to the thermal speed $\sqrt{k_{B}T/M}\sim 4.4$ m/s at 2.39 mK (the minimum for the ideal optical molasses). Interestingly, the $1M$ temperature is less than the ideal case by $\sim 0.9$ mK for larger magnitude detunings, $\Delta_{0}<-\Gamma$.
@@ IV Monte Carlo: antihydrogen trap
Some of the features specific for the optical molasses cooling include the path of the laser which is a straight line with a $2.3^{\circ}$ tilt relative to the trap axis taken to be the $z$-direction. Thus, the $\vec{k}$ is a constant. The motion of the $\bar{\rm H}$ through the trap mixes the different velocity components possibly allowing for three dimensional cooling. Since this mixing is relatively slow, the detuning and intensity need to be chosen to limit the heating of the radial motion due the photon emission step. The beam has a waist of 3.48 mm. On the trap axis, $B_{\perp}/B_{z}<0.0003$ while at a radius of 5 mm the $B_{\perp}/B_{z}<0.05$ which means the direction of $\vec{B}$ has little effect. The laser is pulsed with a duration of 10’s of nanoseconds and with repetition rate of 50 Hz. The laser cooling happens over a time scale measured in hours which is possible due to the cryogenic vacuum.
@@
Figure 3: The magnetic fields for the four simulations described in the text. The red solid line is the “flat” $B$-field, Sec. IV.1 and the blue long dash line is the “harmonic” $B$-field, Sec. IV.3. The green dashed and black dot curves are the slightly dipped and raised traps of Secs. IV.2. The $z$ is the distance from the trap center along the trap axis.
@@
The laser line width, estimated as 60 MHz full-width half-maximum (FWHM), also modestly affects the results since the intrinsic linewidth is approximately 100 MHz. The laser line width is incorporated into the simulation by having each pulse at a slightly different detuning randomly chosen from a Gaussian distribution with a FWHM of 60 MHz. This leads to an effectively broader transition although the effect is not large. We verified that this is a very good approximation (errors less than a couple percent) by comparing to the results from numerically solving the Optical Bloch equations for the same parameters. In all of the figures and results, the linewidth, $\Gamma$, is the natural radiative linewidth of the 2P state.
@@ V Simultaneous cooling 1Sc and 1Sd
The results from simulations in the trap are shown in Fig. 5(b) as the green dash ($\times$) for 2 nJ pulses and maroon dash-dot-dot-dot (solid square) for 1 nJ pulses. There appears to be a small range of detuning (roughly $-3\Gamma<-\Delta_{0}<-2\Gamma$) that leads to decent cooling even with the heating from the blue detuned frequency. This range is narrow because the cooling in the trap without the blue detuned pulses is not good for $-2\Gamma<\Delta_{0}$ and the blue pulses lead to heating for $\Delta_{0}<-3\Gamma$. The two detunings $-2.5$ and $-2\Gamma$ have nearly the same temperature as the one frequency case (these are the black dot (*) and red solid ($+$) curves). The simulations are somewhat unrealistic in that the frequency was held constant for 1 second before switching to a different frequency. In practice, the frequencies probably can not be switched that often. However, it is only necessary to change the frequencies on a time scale where a small number of photons are scattered before the change. Depending on the trap and laser details, probably this would be on the several minute scale. These simulations suggest that the 1Sc and 1Sd states can be simultaneously cooled in the ALPHA trap.

### v3

@@ I Introduction
The most promising cooling technique, and the only one successfully implemented,Baker et al (ALPHA collaboration) is a simple optical molasses based on the 1S-2P transition. The laser in this experiment is a nearly Fourier transform limited pulsed laser with the pulse duration a couple 10’s of nanoseconds and the linewidth of a few 10’s of MHz. For this case, a single 121.6 nm beam, somewhat red detuned from the transition in the magnetic field, cools the axial motion of the $\bar{\rm H}$ atoms. For this geometry, the random re-emission of the photon tends to heat the radial motion. Because the magnetic trapping potentials are not perfectly symmetric, the $\bar{\rm H}$ motion mixes the axial and radial degrees of freedomSurkov et al. (1994); Zhong et al. (2018) which can lead to cooling of all 3 directions.Donnan et al. (2013); Baker et al (ALPHA collaboration) The intensity of the laser has to be sufficiently low that there is time for the mixing to take place between successive photon scattering otherwise the atoms will heat on average. This heating is strongest for smallest detuning where the photon scattering rate is largest. The spatially varying magnetic field also complicates the photoabsorption by giving substantial changes in the detuning versus position in the trap.
@@ II Energy levels in magnetic field
In the ALPHA experiment, $\bar{\rm H}$s are laser cooled in a magnetic trap with a minimum $B\sim 1$ TBaker et al (ALPHA collaboration). This means the energy levels are strongly changed from their low (or zero) field character, Fig. 1. Reference Rasmussen et al. (2017) discusses the energy levels. The energies in a magnetic field are labeled with a Roman letter at the end which, by convention, increases from low to high energy for 1S states while increasing from high to low energy for the 2P states (as ordered for small $B$). There are two 1S states that can be trapped called the 1Sc and 1Sd states. The 1Sd state has the positron and antiproton spin aligned; in the 1Sc state, they are antiparallel. To prevent losses from the photon emission step, the laser will excite the 2Pa state which has the positron spin and orbital angular momentum aligned to give total positron angular momentum $J=3/2,M=3/2$. Because of the different magnetic moments, the transition frequency for the 1Sc-2Pa transition is approximately 675 MHz higher than that for the 1Sd-2Pa transition at 1 T. The difference in frequencies does not change much with increasing $B$; for example, at $B=$0.5 T, the difference in frequencies is 660 MHz while the difference in frequencies is 680 MHz at 1.5 T.
@@
Figure 1: The splittings of the 1S and 2P states relative to their zero field average. The hyperfine terms of the 2P states are too small to resolve. The inset shows the 1Sd and 1Sc states near 1 T.
@@ III.2 Ideal Optical Molasses
In this section, we describe a method for Monte Carlo simulations of laser cooling of $\bar{\rm H}$ using an ideal optical molassesFoot (2005). We use a simplified method that can easily be extended to simulate cooling in the full trap, Sec. IV, but does not include the shift in detuning nor the changing velocity as the $\bar{\rm H}$ moves through the trap but only changes the direction of the velocity vector on a time scale short compared to the times between the laser pulses. Ideal optical molasses cooling assumes the photon has a wave vector, $\vec{k}$, in a random direction relative to the atom’s velocity, $\vec{v}$. The Doppler effect leads to an effective detuning of
@@
Our Monte Carlo simulation of the ideal optical molasses used the following algorithm: (1) for each of $N$ atoms, randomize the direction of the atom’s velocity by keeping its speed but changing the direction by randomly picking a point on the surface of a sphere and for that atom compute the effective detuning, Eq. (4); (2) compute the probability that the photon was absorbed using $P=P_{0}/[1+(2\Delta/\Gamma)^{2}]$ with $P_{0}$ a small number of order 0.01; (3) compute a random number; if it is smaller than $P$, then a photon was absorbed; (4) if a photon is absorbed, compute a random emission direction and use Eq. (5) to update that atom’s velocity. If we wanted to simulate the effect of thermalization from elastic collisions between the atoms, we added a step: (5) randomly pick pairs of atoms and rotate their relative velocity to emulate a collision; repeated often enough step (5) leads to a Maxwell-Boltzmann velocity distribution at temperature $k_{B}T=(2/3)\langle KE\rangle$.
@@
The minimum temperature that can be reached for $1M$ happens for a detuning of $\Delta_{0}\simeq-1.1\Gamma$ and gives $T\simeq 3.05$ mK. The temperature for $1M$ at $\Delta_{0}=-\Gamma/2$ (the minimum of Eq. (6)) is $\simeq 19$ mK which is almost 8×8$\times$8 × hotter than expected from the usual optical molasses relations. From Fig. 2(a), the higher masses become progressively closer to the ideal optical molasses case, red solid line. The reason for the discrepancy is the large size of the velocity kick, 3.26 m/s, compared to the thermal speed $\sqrt{k_{B}T/M}\sim 4.4$ m/s at 2.39 mK (the minimum for the ideal optical molasses). This is similar to the finding in Ref. Castin et al. (1989) where the best detuning and lowest temperature changes as the recoil energy becomes larger than the line width. Interestingly, the $1M$ temperature is less than the ideal case by $\sim 0.9$ mK for larger magnitude detunings, $\Delta_{0}<-\Gamma$.
@@ IV Monte Carlo: antihydrogen trap
Some of the features specific for the optical molasses cooling include the path of the laser which is a straight line with a $2.3^{\circ}$ tilt relative to the trap axis taken to be the $z$-direction. Thus, the $\vec{k}$ is a constant. The motion of the $\bar{\rm H}$ through the trap mixes the different velocity components possibly allowing for three dimensional cooling. Since this mixing is relatively slow, the detuning and intensity need to be chosen to limit the heating of the radial motion due the photon emission step. The beam has a waist of 3.48 mm. On the trap axis, $B_{\perp}/B_{z}<0.0003$ while at a radius of 5 mm the $B_{\perp}/B_{z}<0.05$ which means the direction of $\vec{B}$ has little effect. The laser is pulsed with a duration of 10’s of nanoseconds and with repetition rate of 50 Hz. The intensity is low enough that much less than one photon is scattered from an $\bar{\rm H}$ per pulse on average. The laser cooling happens over a time scale measured in hours which is possible due to the cryogenic vacuum.
@@
Figure 3: The magnetic fields for the four simulations described in the text. The red solid line is the “flat” $B$-field, Sec. IV.1 and the blue long dash line is the “harmonic” $B$-field, Sec. IV.3. The green dashed and black dot curves are the slightly dipped and raised traps of Secs. IV.2. The $z$ is the distance from the trap center along the trap axis. The change in potential energy for the $\bar{\rm H}$ motion is approximately $2/3$ K/T, meaning 0.03 T corresponds to 20 mK $k_{B}$; the change in detuning of the 1Sc or 1Sd to 2Pa is approximately -14 MHz/T meaning 0.01 T change corresponds to 140 MHz redder detuning.
@@
The laser line width, estimated as 60 MHz full-width half-maximum (FWHM), also modestly affects the results since the intrinsic linewidth is approximately 100 MHz. The laser line width is incorporated into the simulation by having each pulse at a slightly different detuning randomly chosen from a Gaussian distribution with a FWHM of 60 MHz. This leads to an effectively broader transition although the effect is not large. This should be a good approximation because the internal states are effectively a two level system (the next allowed transition with similar transition matrix element is detuned by approximately 300 linewidths), the laser pulses are separated by sufficient time that there are no atomic coherences (either in position or internal states) between the pulses, the laser pulses are nearly Fourier transform limited with a duration more than 10 times longer than the lifetime, and the peak Rabi frequency is much less than the 2P decay rate. We verified that this is a very good approximation (errors less than a couple percent) by comparing to the results from numerically solving the Optical Bloch equations for the same laser parameters. In all of the figures and results, the linewidth, $\Gamma$, is the natural radiative linewidth of the 2P state.
@@ V Simultaneous cooling 1Sc and 1Sd
The results from simulations in the trap are shown in Fig. 5(b) as the green dash ($\times$) for 2 nJ pulses and maroon dash-dot-dot-dot (solid square) for 1 nJ pulses. There appears to be a small range of detuning (roughly $-3\Gamma<\Delta_{0}<-2\Gamma$) that leads to decent cooling even with the heating from the blue detuned frequency. This range is narrow because the cooling in the trap without the blue detuned pulses is not good for $-2\Gamma<\Delta_{0}$ and the blue pulses lead to heating for $\Delta_{0}<-3\Gamma$. The two detunings $-2.5$ and $-2\Gamma$ have nearly the same temperature as the one frequency case (these are the black dot (*) and red solid ($+$) curves). The simulations are somewhat unrealistic in that the frequency was held constant for 1 second before switching to a different frequency. In practice, the frequencies probably can not be switched that often. However, it is only necessary to change the frequencies on a time scale where a small number of photons are scattered before the change. Depending on the trap and laser details, probably this would be on the several minute scale. These simulations suggest that the 1Sc and 1Sd states can be simultaneously cooled in the ALPHA trap.